CONSTANT GROWTH VALUATION

Holtzman Clothiers's stock currently sells for $38 a share. It just paid a dividend of $3.5 a share (i.e., D0 = $3.5). The dividend is expected to grow at a constant rate of 9% a year.

What stock price is expected 1 year from now? Round your answer to two decimal places.

Please state the formulas clearly to help me understand.


$______

What is the required rate of return? Round your answer to two decimal places. Do not round your intermediate calculations.
_______ %
Answer #1
1. Computation of Price of Share on eyear from now
Po=$38
D0=$3.5
g=9%
r= 19%
D1= Do(1+g)= $3.5(1.09)=$3.815
P1=D1(1+g)/R-g
=$3.815(1.09)/19%-9%
= $41.58
2. Computation of required reate of return
Po=$38
D0=$3.5
g=9%
r= Required rate of Return
R= D0(1+g)/P0 +g
R= $3.5(1.09)/$38+0.09